OpenStudy (anonymous):

Write the equation of the line with a slope of -1/3 and passing through the point (6, -4).

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

wait im confused to what that is?

OpenStudy (anonymous):

m is the slope and x1,y1 the point u have

OpenStudy (anonymous):

im not exactly understanding..."/

OpenStudy (anonymous):

what is the slope?

OpenStudy (anonymous):

-1/3

OpenStudy (anonymous):

ok plug that into y-y1=m(x-x1) now what is the x and y values?

OpenStudy (anonymous):

0,3?

OpenStudy (mathlegend):

In that formula.. m = slope

OpenStudy (mathlegend):

@itsjustme_lol can you plug in everything into that formula now?

OpenStudy (anonymous):

so m=-1/3

OpenStudy (mathlegend):

Right.... and you know what x1 and y1 are since they gave you the coordinates. So plug it in.... y-y1=m(x-x1)

OpenStudy (anonymous):

y-4=m(x-6)?

OpenStudy (mathlegend):

y-(-4)=-1/3(x-6) y+4=-1/3(x-6) @itsjustme_lol do you understand that?

OpenStudy (anonymous):

why did we put +4 in the second one?

OpenStudy (mathlegend):

y-(-4) = y+4

OpenStudy (mathlegend):

two negatives makes a positive

OpenStudy (mathlegend):

It is a negative being multiplied by a negative. That is something you should be familiar with.

OpenStudy (mathlegend):

y+4=-1/3(x-6) Can you simplify the right side of this equation?

OpenStudy (anonymous):

give me a second

OpenStudy (anonymous):

wait im still confused. im sorry!!!

OpenStudy (mathlegend):

It is fine.

OpenStudy (mathlegend):

What are you confused about?

OpenStudy (anonymous):

i dont know how to solve that equation

OpenStudy (mathlegend):

y+4=-1/3(x-6) *We must first distribute that -1/3 on the right side of the equation. y+4=-1/3x+2 Now we need to make this equation into the slope-intercept form... so y=mx+b... so lets subtract the 4 from both sides. y+4=-1/3x+2 -4 -4 --------------------- y=-1/3x-2

OpenStudy (mathlegend):

@itsjustme_lol do you understand?

OpenStudy (anonymous):

that makes sense now